OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <6, -2>, v = <8, 24>

OpenStudy (anonymous):

if they are perpendicular, the the dot product is zero

OpenStudy (anonymous):

which i think it i in this case as \[6\times 8+(-2)\times 24=0\]

OpenStudy (anonymous):

A salesman drives from Ajax to Barrington, a distance of 120 mi, at a steady speed. He then increases his speed by 10 mi/h to drive the 150 mi from Barrington to Collins. If the second leg of his trip took 6 min more time than the first leg, how fast was he driving between Ajax and Barrington? Omg sorry but please help me I've been on here for a while!

OpenStudy (anonymous):

@Alohi27 repost and i will see if i can help with this problem

OpenStudy (anonymous):

I did...

OpenStudy (anonymous):

umm i need heppp

OpenStudy (anonymous):

helpp

OpenStudy (anonymous):

ill solve it and give my answer

OpenStudy (anonymous):

i answered above take the dot product, see that you get zero that means they are perpendicular

OpenStudy (anonymous):

oh okay thank u!